OpenStudy (anonymous):

find the zeroes and state the multiplicity of multiple zeroes y=(x+1)^2 (x-1) (x-2)

jhonyy9 (jhonyy9):

do you know when will be this equation equal zero ?

jhonyy9 (jhonyy9):

so when one from these factors will be equal zero so do you know how you get the zeroes of these factors ?

OpenStudy (anonymous):

no

jhonyy9 (jhonyy9):

so in this way --- separated make every parentheses equal zero : (x+1)^2 = 0 -and solve it for x same like the first x-1=0 and x-2=0

jhonyy9 (jhonyy9):

can you solve it ?

OpenStudy (anonymous):

x=1 x=1 x=2

OpenStudy (anonymous):

x=-1

jhonyy9 (jhonyy9):

no step by step 1. (x+1)=0

jhonyy9 (jhonyy9):

sorry (x+1)^2 =0

jhonyy9 (jhonyy9):

courage - this is easy

OpenStudy (anonymous):

is it not -1

jhonyy9 (jhonyy9):

so (x+1)^2 =0 yes x=-1 is right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i find the multiplicities

jhonyy9 (jhonyy9):

but there are again two zeroes what you need calculi again

jhonyy9 (jhonyy9):

x=-1 is the first one you need getting again two

OpenStudy (anonymous):

getting again two?

jhonyy9 (jhonyy9):

just make equal zero the second and 3rd parentheses and calculi for x

jhonyy9 (jhonyy9):

exactly how you have solved the first

OpenStudy (anonymous):

x=1 x=2

jhonyy9 (jhonyy9):

x-1=0 x=?

OpenStudy (anonymous):

1

jhonyy9 (jhonyy9):

yes 1 and 2 right sure

jhonyy9 (jhonyy9):

so than now what are zeroes of this equation ?

OpenStudy (anonymous):

-1, 1, 2

jhonyy9 (jhonyy9):

yes - and what mean multiplicity of them ?

OpenStudy (anonymous):

i dont know what that means